OpenStudy (anonymous):

Find the 6th term of a geometric sequence t3 = 444 and t7 = 7104

OpenStudy (anonymous):

use common ratio t7/t6=a t6/t5=a . . . t4/t3=a t7=t3*\[t7=t3*a ^{4}\] t7/t3=a^4 thus a=2

OpenStudy (anonymous):

thus t6=t7/a=t7/2=3552 t5=t6/2=1776 t4=t5/2=888 t3=t4/2=444 t2=t3/2=222 t1=t2/2=111